OpenStudy (anonymous):

Will Fan and Medal! Plz work out problem below. A train traveled 210 miles in 3.5 hours to its first stop. It then continued on traveling another 130 miles in 2.5 hours. What was the average speed of the train for the entire journey? A. 52 mi/h B. 56 mi/h C. 56.7 mi/h D. 60 mi/h

OpenStudy (anonymous):

@GracieBugg @DarkMoonZ @anyone @bohotness @Bookworm14

OpenStudy (bohotness):

lol i am here :P

OpenStudy (anonymous):

Can you help?

OpenStudy (bohotness):

Divide total miles by total hours.

OpenStudy (anonymous):

I have on both and I come with 60 and 52.

OpenStudy (bohotness):

what is the toals of hours?

OpenStudy (anonymous):

Not sure that is waht I came up with

OpenStudy (anonymous):

@Luigi0210

OpenStudy (bohotness):

3.5 hours + 2.5 hours. =?

OpenStudy (bookworm14):

So for this you are needing to find the unit rate or in other words, how many miles the train travels in 1 hour. In order to get this answer you must add the total miles together. Then you must add the amounts of time together. 130+210 =? 3.5+2.5= ?

OpenStudy (anonymous):

6

OpenStudy (bookworm14):

3.5+2.5 = 6 hours now what is 210 + 130?

OpenStudy (anonymous):

340

OpenStudy (bookworm14):

now you have to take the total miles and divide it by the total hours. 340/6= ???

OpenStudy (anonymous):

56.66

OpenStudy (bookworm14):

So we would round that to 56.7 I am not sure if they are asking for and B or C.... I would go with C, because they say what is the average speed, they DO NOT ask what is the "approximate" average speed

OpenStudy (bookworm14):

did that help any? or are you still confused? If need be, I can work through an example with you